The number of cinema tickets in Karachi and the ticket fares are regulated. The fare currently charged is Rs.500 a ticket. The Cinema Association want to obtain government's permission to raise the fare to increase their revenues and ask you to be their economic adviser. After studying the market, you come up with the following demand schedule for cinema tickets:


Price (Rupees per ticket) Quantity demanded (tickets per month)

300 160

400 120

500 80

600 40

700 0


a) Calculate the price elasticity of demand for cinema tickets as the fare rises from Rs.500 to Rs.600. Is the demand price elastic or inelastic for this fare rise?


1
Expert's answer
2020-12-07T07:15:40-0500

"Price elasticity=\frac{Percentage Change in Quantity demanded}{Percentage change in price}"  

When fare price rises from RS 500 to RS 600, the number of tickets demanded falls to 40 from 80.

%change in price "\frac{(600-500)}{500}\times100=20%"%


%change in Quantity Demanded "\frac{(40-80)}{80}\times100=-50"%


Therefore, PED"=\frac{50}{-20}=-2.5"

In this case, demand is price elastic since the "PED>1"